# Fortify the Guardian, Not the Treasure: Resilient Adversarial Detectors

### v1

@@ Abstract
This paper presents RADARRobust Adversarial Detection via Adversarial Retraining—an approach designed to enhance the robustness of adversarial detectors against adaptive attacks, while maintaining classifier performance. An adaptive attack is one where the attacker is aware of the defenses and adapts their strategy accordingly. Our proposed method leverages adversarial training to reinforce the ability to detect attacks, without compromising clean accuracy. During the training phase, we integrate into the dataset adversarial examples, which were optimized to fool both the classifier and the adversarial detector, enabling the adversarial detector to learn and adapt to potential attack scenarios. Experimental evaluations on the CIFAR-10 and SVHN datasets demonstrate that our proposed algorithm significantly improves a detector’s ability to accurately identify adaptive adversarial attacks—without sacrificing clean accuracy.
@@ 1 Introduction
Rapid advances in deep learning systems and their applications in various fields such as finance, healthcare, and transportation, have greatly enhanced human capabilities [1, 2, 3, 4, 5]. However, these systems continue to be vulnerable to adversarial attacks, where unintentional or intentionally designed inputs—known as adversarial examples—can mislead the decision-making process [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Such adversarial attacks pose severe threats both to the usage and trustworthiness of deep learning technologies in critical applications [16, 17].
Extensive research efforts have been dedicated to developing robust machine learning classifiers that are resilient to adversarial attacks. One popular and effective way relies on adversarial training, which involves integrating adversarial examples within the training process, to improve the classifier’s resistance to attack [18, 19]. However, the emergence of adversarial detectors designed to identify malicious inputs [20, 21, 22, 23, 24] has opened new avenues for enhancing or thwarting the security of AI systems.
@@
Several studies have found that these detectors—even when combined with robust classifiers that have undergone adversarial training—can be fooled by designing tailored attacks; this engenders a cat-and-mouse cycle of attacking and defending [25, 26]. It is therefore critical to thoroughly assess the robustness of adversarial detectors and ensure that they are able to withstand adaptive adversarial attacks.
@@
  • Paradigm shift: Instead of robustifying the classifier, we robustify the adversarial detector.
  • We propose RADAR, a novel adversarial training approach to improve the resilience of adversarial detectors.
  • Our study conducts extensive evaluations on two datasets and six detection architectures, providing insights into the generalizability and efficacy of adversarial training across varied data distributions and model architectures.
@@ 2 Previous Work
The foundation of adversarial training was laid by Madry et al. [18], who demonstrated its effectiveness on the MNIST [27] and CIFAR-10 [28] datasets. Their approach used Projected Gradient Descent (PGD) attacks to generate adversarial examples and incorporate them into the training process alongside clean (unattacked) data.
@@
Subsequent work by Kurakin et al. [29] explored diverse adversarial training methods, highlighting the trade-off between robustness and accuracy. These early pioneering works laid the groundwork for the field of adversarial training and showcased its potential for improving classifier robustness.
@@
Adversarial training has evolved significantly since its inception. Zhang et al. [30] introduced MixUp training, probabilistically blending clean and adversarial samples during training, achieving robustness without significant accuracy loss.
@@
Tramer and Boneh [31] investigated multi-attack adversarial training, aiming for broader robustness against diverse attack types.
@@
A number of works explored formal-verification frameworks to mathematically certify classifier robustness against specific attacks, demonstrating provably robust defense mechanisms [32, 33, 34, 35].
@@
Zhang et al. [36] provided theoretical insights into the robustness of adversarially trained models, offering a deeper understanding of the underlying mechanisms and limitations. Their work elucidated the trade-offs between robustness, expressiveness, and generalization in adversarial training paradigms, contributing to a more comprehensive theoretical framework for analyzing and designing robust classifiers.
@@
In the pursuit of improving robustness against adversarial examples, Xie et al. [37] proposed a novel approach using feature denoising. Their work identified noise within the extracted features as a vulnerability during adversarial training. To address this, they incorporated denoising blocks within the network architecture of a convolutional neural network (CNN). These blocks leverage techniques such as non-local means filters to remove the adversarial noise, leading to cleaner and more-robust features. Their approach achieved significant improvements in adversarial robustness on benchmark datasets, when compared to baseline adversarially trained models. This work highlights feature denoising as a promising defense mechanism that complements existing adversarial training techniques.
@@
Building on the limitations of standard adversarial training, requiring a large amount of labeled data, Carmon et al. [38] explored leveraging unlabeled data for improved robustness. Their work demonstrates that incorporating unlabeled data through semi-supervised learning techniques, such as self-training, can significantly enhance adversarial robustness. This approach bridges the gap in sample complexity, achieving high robust accuracy with the same amount of labeled data needed for standard accuracy. Their findings were validated empirically on datasets such as CIFAR-10, achieving state-of-the-art robustness against various adversarial attacks. This research highlights the potential of semi-supervised learning as a cost-effective strategy to improve adversarial training and achieve robust models.
@@
Despite its successes, adversarial training faces several challenges. Transferability of adversarial examples remains a concern [39, 40], because attacks often lack effectiveness across different classifiers and architectures. The computational cost of generating and training on adversarial examples can be significant, especially for large models and datasets. Moreover, adversarial training can lead to a reduction in clean-data accuracy, requiring effective optimization strategies to mitigate this trade-off [41]. Additionally, carefully crafted evasive attacks can sometimes bypass adversarially trained classifiers, highlighting the need for further research and diversification of defense mechanisms.
@@
There are but a few examples of works that compare adversarial attacks that take into account full knowledge of the defense, i.e., adaptive attacks. He et al. [42] showed promising results, at the cost of multiple model inferences coupled with multiple augmented neighborhood images. Yang et al. [43] presented a novel encoder-generator architecture that compared the generated image label with the original image label. We believe this was a step in the right direction, yet with a computationally expensive generator that had difficulties differentiating classes that are semantically close. Klingner et al. [44] demonstrated a detection method based on edge extraction of features such as depth estimation and semantic segmentation, and comparison to natural image edges based on the SSIM metric. We believe the underlying assumption of adversarial examples containing abnormal edges might not hold in settings that are not natural.
@@ 3.1 Threat Model
$\operatorname*{arg\,max}_{k=0,1,...K-1}f_{k}(x+\delta)\neq y,\;(x,y)\in\mathcal{D},\;\;\text{s.t.}\;\|\delta\|_{p}\leq\epsilon.$ (1)
@@ 3.2 Problem Definition
But a potential conflict manifests with this objective. While gradient descent allows for simultaneous maximization of the cross-entropy loss ($\mathcal{L}_{\text{CE}}$) and the binary cross-entropy loss ($\mathcal{L}_{\text{BCE}}$), they might contradict each other, adopting different optimization strategies. To address this challenge we employ the selective and orthogonal approaches proposed Bryniarski et al. [45].
Selective Projected Gradient Descent (SPGD) [45] optimizes only with respect to a constraint that has not been satisfied yet—while not optimizing against a constraint that has been optimized. The loss function used by SPGD is:
@@
Orthogonal Projected Gradient Descent (OPGD) [45] focuses on modifying the update direction to ensure it remains orthogonal to the constraints imposed by previously satisfied objectives. This orthogonal projection ensures that the update steers the input towards the adversarial objective without violating the constraints imposed by the classifier’s decision boundary, allowing for the creation of adversarial examples that bypass the detector. Thus, the update rule is slightly different:
@@ 4 Experimental Framework
Datasets and models. We used the VGG-{11,13,16} and ResNet-{18, 34, 50} architectures both for classification and for adversarial detection. We modified the classification head of the detector, $C\in\mathbb{R}^{K\times e}$, where $K$ is the number of classes and $e$ is the embedding-space size, to $C\in\mathbb{R}^{2\times e}$, for binary classification (benign/adversarial), and added a sigmoid transformation at the end.
We used the definition of attack success rate presented by Bryniarski et al. [45] as part of their evaluation methodology. Attack efficacy is measured through Attack Success Rate at N, SR@N: proportion of deliberate attacks achieving their objectives subject to the condition that the defense’s false-positive rate is configured to N%. The underlying motivation is that in real-life scenarios we must strike a delicate balance between security and precision. A 5% false positive is acceptable, while extreme cases might even use 50%. Our results proved excellent so we set N to a low 5%, i.e., SR@5.
Throughout the experiments, the allowed perturbation was constrained by an L-infinity norm, denoted as $\|\cdot\|_{\infty}$, with a maximum magnitude set to $\epsilon=16/255$. We employed an adversarial attack strategy, specifically utilizing 100 iterations of PGD with a step size parameter $\alpha$ set to 0.03. This approach was employed to generate adversarial instances and assess the resilience of the classifier under scrutiny. We then evaluated the classifiers on the attacked test set—the results are delineated in Table 1.
@@
Dataset Classifier Accuracy (ben) Accuracy (adv)
CIFAR-10 VGG-11 92.40 0.35
VGG-13 94.21 0.19
VGG-16 94.00 5.95
ResNet-18 92.60 0.14
ResNet-34 93.03 0.21
ResNet-50 93.43 0.26
SVHN VGG-11 93.96 0.00
VGG-13 94.85 0.01
VGG-16 94.95 0.48
ResNet-18 94.88 0.02
ResNet-34 94.84 0.11
ResNet-50 94.33 0.10
Afterwards, we split the training data into training ($70\%$) and validation sets ($30\%$). We trained 3 VGG-based and 3 ResNet-based adversarial detectors for $20$ epochs, using the Adam optimizer (with default $\beta_{1}$ and $\beta_{2}$ values), with a learning rate of $41e-$ 4, CosineAnnealing learning-rate scheduler with $T_{\text{max}}=10$, and a batch size of $256$. We tested the adversarial detectors on the test set that was comprised of one half clean images and one half attacked images: They all performed almost perfectly.
After the initial clean training phase we performed adversarial finetuning on the adversarial detectors. This involved attacking them with an adaptive adversarial attack, e.g., OPGD. We ran the adversarial finetuning for $20$ epochs, using the same optimizer, ReduceLROnPlateau learning rate scheduler with patience parameter set to $3$, and batch size of $256$. Then, we tested on a test set—with equal distribution of clean images and OPGD-attacked images.

### v2

@@ Abstract
This paper presents RADARRobust Adversarial Detection via Adversarial Retraining—an approach designed to enhance the robustness of adversarial detectors against adaptive attacks, while maintaining classifier performance. An adaptive attack is one where the attacker is aware of the defenses and adapts their strategy accordingly. Our proposed method leverages adversarial training to reinforce the ability to detect attacks, without compromising clean accuracy. During the training phase, we integrate into the dataset adversarial examples, which were optimized to fool both the classifier and the adversarial detector, enabling the adversarial detector to learn and adapt to potential attack scenarios. Experimental evaluations on CIFAR-10, SVHN and ImageNet datasets demonstrate that our proposed algorithm significantly improves a detector’s ability to accurately identify adaptive adversarial attacks—without sacrificing clean accuracy.
@@ 1 Introduction
Rapid advances in deep learning systems and their applications in various fields such as finance, healthcare, and transportation, have greatly enhanced human capabilities (Huang et al., 2020; Miotto et al., 2018; Chen et al., 2024a; Shamshirband et al., 2021; Lv et al., 2020). However, these systems continue to be vulnerable to adversarial attacks, where unintentional or intentionally designed inputs—known as adversarial examples—can mislead the decision-making process (Goodfellow et al., 2014; Szegedy et al., 2013; Tamam et al., 2023; Chen et al., 2024b; Lapid et al., 2024; Carlini & Wagner, 2017b; Lapid & Sipper, 2023b; Andriushchenko et al., 2020; Lapid et al., 2022; Lapid & Sipper, 2023a). Such adversarial attacks pose severe threats both to the usage and trustworthiness of deep learning technologies in critical applications (Finlayson et al., 2019; Liu et al., 2023).
Extensive research efforts have been dedicated to developing robust machine learning classifiers that are resilient to adversarial attacks. One popular and effective way relies on adversarial training, which involves integrating adversarial examples within the training process, to improve the classifier’s resistance to attack (Madry et al., 2017; Zhao et al., 2022). However, the emergence of adversarial detectors designed to identify malicious inputs (Lu et al., 2017; Grosse et al., 2017a; Gong & Wang, 2023; Lust & Condurache, 2020; Metzen et al., 2016) has opened new avenues for enhancing or thwarting the security of AI systems.
@@
Several studies have found that these detectors—even when combined with robust classifiers that have undergone adversarial training—can be fooled by designing tailored attacks; this engenders a cat-and-mouse cycle of attacking and defending (Carlini & Wagner, 2017a; Grosse et al., 2017b). It is therefore critical to thoroughly assess the robustness of adversarial detectors and ensure that they are able to withstand adaptive adversarial attacks.
@@
  • Paradigm shift: Unlike existing approaches that focus on robustifying classifiers, we introduce a novel paradigm of robustifying adversarial detectors, eliminating the trade-off between clean and robust classification.
  • We introduce RADAR, a pioneering adversarial training technique specifically designed to enhance the resilience of adversarial detectors, demonstrating its effectiveness through extensive evaluations on three datasets and across six detection architectures.
  • Our study provides comprehensive insights into the generalizability and efficacy of adversarial training strategies across diverse data distributions and model architectures, reinforcing the critical role of robust adversarial detectors in securing machine learning systems.
@@ 2 Previous Work
The foundation of adversarial training was laid by Madry et al. (2017), who demonstrated its effectiveness on the MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky, 2009) datasets. Their approach used Projected Gradient Descent (PGD) attacks to generate adversarial examples and incorporate them into the training process alongside clean (unattacked) data.
@@
Subsequent work by Kurakin et al. (2016) explored diverse adversarial training methods, highlighting the trade-off between robustness and accuracy. These early pioneering works laid the groundwork for the field of adversarial training and showcased its potential for improving classifier robustness.
@@
Adversarial training has evolved significantly since its inception. Zhang et al. (2017) introduced MixUp training, probabilistically blending clean and adversarial samples during training, achieving robustness without significant accuracy loss.
@@
Tramer & Boneh (2019) investigated multi-attack adversarial training, aiming for broader robustness against diverse attack types.
@@
A number of works explored formal-verification frameworks to mathematically certify classifier robustness against specific attacks, demonstrating provably robust defense mechanisms (Cohen et al., 2019; Lecuyer et al., 2019; Li et al., 2023; Singh et al., 2018).
@@
Zhang et al. (2019) provided theoretical insights into the robustness of adversarially trained models, offering a deeper understanding of the underlying mechanisms and limitations. Their work elucidated the trade-offs between robustness, expressiveness, and generalization in adversarial training paradigms, contributing to a more comprehensive theoretical framework for analyzing and designing robust classifiers.
@@
In the pursuit of improving robustness against adversarial examples, Xie et al. (2019) proposed a novel approach using feature denoising. Their work identified noise within the extracted features as a vulnerability during adversarial training. To address this, they incorporated denoising blocks within the network architecture of a convolutional neural network (CNN). These blocks leverage techniques such as non-local means filters to remove the adversarial noise, leading to cleaner and more-robust features. Their approach achieved significant improvements in adversarial robustness on benchmark datasets, when compared to baseline adversarially trained models. This work highlights feature denoising as a promising defense mechanism that complements existing adversarial training techniques. Pinhasov et al. (2024) introduced a novel method to combat adversarial attacks on deepfake detectors. The authors leveraged eXplainable Artificial Intelligence (XAI) (Došilović et al., 2018) to generate interpretability maps, revealing the decision-making process of the deepfake detector. By analyzing these maps, their system could identify vulnerabilities exploited by adversarial attacks.
@@
Building on the limitations of standard adversarial training, requiring a large amount of labeled data, Carmon et al. (2019) explored leveraging unlabeled data for improved robustness. Their work demonstrates that incorporating unlabeled data through semi-supervised learning techniques, such as self-training, can significantly enhance adversarial robustness. This approach bridges the gap in sample complexity, achieving high robust accuracy with the same amount of labeled data needed for standard accuracy. Their findings were validated empirically on datasets such as CIFAR-10, achieving state-of-the-art robustness against various adversarial attacks. This research highlights the potential of semi-supervised learning as a cost-effective strategy to improve adversarial training and achieve robust models.
@@
Despite its successes, adversarial training faces several challenges. Transferability of adversarial examples remains a concern (Cheng et al., 2019; Dong et al., 2021), because attacks often lack effectiveness across different classifiers and architectures. The computational cost of generating and training on adversarial examples can be significant, especially for large models and datasets. Moreover, adversarial training can lead to a reduction in clean-data accuracy, requiring effective optimization strategies to mitigate this trade-off (Yang et al., 2020). Additionally, carefully crafted evasive attacks can sometimes bypass adversarially trained classifiers, highlighting the need for further research and diversification of defense mechanisms.
@@
There are but a few examples of works that compare adversarial attacks that take into account full knowledge of the defense, i.e., adaptive attacks. He et al. (2022) showed promising results, at the cost of multiple model inferences coupled with multiple augmented neighborhood images. Klingner et al. (2022) demonstrated a detection method based on edge extraction of features such as depth estimation and semantic segmentation, and comparison to natural image edges based on the SSIM metric. We believe the underlying assumption of adversarial examples containing abnormal edges might not hold in unnatural settings.
Previous research has explored using sentiment analysis for adversarial example detection (Wang et al., 2023), focusing on the impact of adversarial perturbations on deep neural networks’ hidden-layer feature maps under attack.
All of the above compared their work with some adaptive power like PGD, yet didn’t show state-of-the-art adaptive attacks, which optimize under the constraint imposed by the classifier decision boundary, like SPGD and OPGD (Bryniarski et al., 2021). Only a few methods were compared under these powerful optimizers. Yang et al. (2022a) presented a novel encoder-generator architecture that compared the generated image label with the original image label. We believe this was a step in the right direction, yet with a computationally expensive generator that had difficulties differentiating semantically close classes. We compared our results to theirs, and others (Shan et al., 2020; Sperl et al., 2020; Tian et al., 2021), which already have some comparison to SPGD or OPGD.
@@ 3.1 Threat Model
$\operatorname*{arg\,max}_{k=0,1,...K-1}f_{k}(x+\delta)\neq y,\;(x,y)\in\mathcal{D},\;\;\text{s.t.}\;\|\delta\|_{p}\leq\epsilon,$ (1)
where $K$ denotes the number of classes in the classification task. We can solve this problem using the following optimization objective:
@@ 3.2 Problem Definition
But a potential conflict manifests with this objective. While gradient descent allows for simultaneous maximization of the cross-entropy loss ($\mathcal{L}_{\text{CE}}$) and the binary cross-entropy loss ($\mathcal{L}_{\text{BCE}}$), they might contradict each other, adopting different optimization strategies. The $L_{CE}$ is designed to improve the classifier’s ability to correctly classify inputs by minimizing the classification error. This objective drives $f_{\theta}$ to adjust its parameters such that the predicted class probabilities align closely with the true class labels. Conversely, the $L_{BCE}$ aims to enhance the adversarial detector’s robustness by distinguishing between clean and adversarial examples. This involves maximizing the detector’s sensitivity to adversarial perturbations, thus encouraging the model, $g_{\phi}$, to identify and separate adversarial inputs from genuine ones.
When these objectives are combined in a regular Lagrange multiplier formulation, the optimization process can encounter inherent contradictions. The classifier’s goal to minimize $L_{CE}$ pushes the model towards a decision boundary that accurately classifies clean examples, which may inadvertently align with the directions that adversarial examples exploit. On the other hand, the detector’s goal to minimize $L_{BCE}$ focuses on altering the decision boundary to detect adversarial perturbations, potentially at the expense of the classifier’s accuracy on clean examples.
This dual optimization can result in non-convergent behavior due to the following reasons. Firstly, the gradients derived from $L_{CE}$ and $L_{BCE}$ may point in different directions. For instance, a gradient step that improves classification accuracy might reduce the detector’s ability to identify adversarial examples, and vice versa. This tug-of-war can hinder the optimization process, preventing either objective from being fully realized. Secondly, the model’s capacity to adjust its parameters is limited. Allocating resources to improve robustness against adversarial attacks might detract from the resources available for improving classification accuracy. This competition for finite model capacity can lead to a suboptimal trade-off between robustness and accuracy. Lastly, the optimization landscape of $L_{CE}$ and $L_{BCE}$ can dynamically change as the model parameters are updated. The evolving nature of adversarial attacks means that the detector’s objective is continually shifting, which can destabilize the classifier’s learning process and lead to oscillatory or divergent behavior. To address this challenge we employ the selective and orthogonal approaches proposed by Bryniarski et al. (2021).
Selective Projected Gradient Descent (SPGD) (Bryniarski et al., 2021) optimizes only with respect to a constraint that has not been satisfied yet—while not optimizing against a constraint that has been optimized. The loss function used by SPGD is:
@@
Orthogonal Projected Gradient Descent (OPGD) (Bryniarski et al., 2021) focuses on modifying the update direction to ensure it remains orthogonal to the constraints imposed by previously satisfied objectives. This orthogonal projection ensures that the update steers the input towards the adversarial objective without violating the constraints imposed by the classifier’s decision boundary, allowing for the creation of adversarial examples that bypass the detector. Thus, the update rule is slightly different:
@@ 4 Experimental Framework
Datasets and models. We used the VGG-{11,13,16} and ResNet-{18, 34, 50} architectures both for classification and for adversarial detection. Specifically, we employed these architectures in dual roles: as classifiers to perform the primary task of image classification and as adversarial detectors to identify adversarial examples. By leveraging the same architectures for both tasks, we ensured consistency in our experimental setup and enabled a comprehensive evaluation of their robustness and effectiveness in the context of adversarial training. We modified the classification head of the detector, $C\in\mathbb{R}^{K\times e}$, where $K$ is the number of classes and $e$ is the embedding-space size, to $C\in\mathbb{R}^{2\times e}$, for binary classification (benign/adversarial), and added a sigmoid transformation at the end.
We utilized the CIFAR-10 (Krizhevsky, 2009), SVHN (Netzer et al., 2011) and a subset of ImageNet (Deng et al., 2009) dataset to evaluate our proposed approach. For ImageNet, we randomly selected 50 classes from the original dataset, which contains over 14 million images and 1,000 classes. This subset was used to maintain manageability and to focus our evaluation on a representative sample of the larger dataset, while still leveraging the diversity and complexity that ImageNet provides for benchmarking in image classification tasks. It is important to highlight that adversarial training for adversarial detectors incurs substantially higher computational costs compared to conventional adversarial training. As a result, this process is considerably more time-consuming when applied to large datasets. To ensure the feasibility of our experiments, we therefore employed a reduced number of classes from ImageNet.
We used the definition of attack success rate presented by Bryniarski et al. (2021) as part of their evaluation methodology. Attack efficacy is measured through Attack Success Rate at N, SR@N: proportion of deliberate attacks achieving their objectives subject to the condition that the defense’s false-positive rate is configured to N%. The underlying motivation is that in real-life scenarios we must strike a delicate balance between security and precision. A 5% false positive is acceptable, while extreme cases might even use 50%. Our results proved excellent so we set N to a low 5%, i.e., SR@5.
Throughout the experiments, the allowed perturbation was constrained by an L-infinity norm, denoted as $\|\cdot\|_{\infty}$, with a maximum magnitude set to $\epsilon=\frac{16}{255}$. We employed an adversarial attack strategy, specifically utilizing 100 iterations of PGD with a step size parameter $\alpha$ set to 0.03. This approach was employed to generate adversarial instances and assess the resilience of the classifier under scrutiny. We then evaluated the classifiers on the attacked test set—the results are delineated in Table 1.
@@
Dataset VGG-11 VGG-13 VGG-16 ResNet-18 ResNet-34 ResNet-50
Ben Adv Ben Adv Ben Adv Ben Adv Ben Adv Ben Adv
CIFAR-10 92.40 0.35 94.21 0.19 94.00 5.95 92.60 0.14 93.03 0.21 93.43 0.26
SVHN 93.96 0.00 94.85 0.01 94.95 0.48 94.88 0.02 94.84 0.11 94.33 0.10
ImageNet 70.08 0.00 70.44 0.00 71.96 0.00 70.24 0.00 72.24 0.00 74.96 0.00
Afterwards, we split the training data into training ($70\%$) and validation sets ($30\%$). We trained 3 VGG-based and 3 ResNet-based adversarial detectors for $20$ epochs, using the Adam optimizer (with default $\beta_{1}$ and $\beta_{2}$ values), with a learning rate of $41e-$ 4, CosineAnnealing learning-rate scheduler with $T_{\text{max}}=10$, and a batch size of $32$. We tested the adversarial detectors on the test set that was comprised of one half clean images and one half attacked images: They all performed almost perfectly.
Following the initial clean training phase, we implemented our proposed RADAR approach, which involves adversarial fine-tuning on the adversarial detectors. This process entailed attacking the detectors to an adaptive adversarial attack, specifically the OPGD attack. The adversarial fine-tuning was conducted over the course of 20 epochs, utilizing the same optimizer and the ReduceLROnPlateau learning rate scheduler, with the patience parameter set to 3, and a batch size of 32. Subsequently, the performance was evaluated on a test set consisting of an equal distribution of clean images and images subjected to OPGD attacks.